what is the sum of first 100 natural numabers

Grade:12

9 Answers

KALYAN
8 years ago
5050
formula is n(n+1) divided by 2
venky
8 years ago
5050
KARTHIK
8 years ago
5050
SAPTAGIRI
8 years ago
5050
Mondo Camry
8 years ago
use n(n+1)/2 and write n as 100 , u will get 100x101/2=5050
sri tanish
8 years ago
5050
 
SHANMUKESHWAR
8 years ago
n(n+1)/2=100*101/2=5050
DURGARAM
8 years ago
5050
4222
8 years ago
5050